OpenStudy (anonymous):

PLEASE HELP ME Solve, using the substitution method. y = 4x + 6 5x – y = 6 There are an infinite number of solutions. There is no solution. (18, 12) (12, 54)

OpenStudy (anonymous):

We have an equation for y, so put this into the second one and we get: 5x-(4x+6)=6 x-6=6 x=12 Put this back into the first equation to get: y=4*12+6=54 The solution is (12, 54) :)

OpenStudy (anonymous):

thanks

OpenStudy (dtan5457):

The answer is 12, 54. y=4x+6 5x-y=6 5x-6=4x+6 Do the subtraction etc and you get x=12. Plug that back to the first equation and you get 54.

OpenStudy (anonymous):

thaanks